OpenStudy (noah690):

f(x)=x^2-3x and g(x)=4x^5 what is f(x) * g(x). Is the answer 4x^7-12x^6 Thanks.

OpenStudy (solomonzelman):

\(\large\color{red}{ \displaystyle f(x)=x^2-3x }\) \(\large\color{blue}{ \displaystyle g(x)=4x^5 }\) \(\large\color{red}{ \displaystyle f(x) }\) \(\large\color{black}{ \times }\) \(\large\color{blue}{ \displaystyle g(x) \color{black}{=} }\)\(\large\color{red}{ \displaystyle \left(x^2-3x\right) }\) \(\large\color{black}{ \times }\) \(\large\color{blue}{ \displaystyle \left(4x^5\right) \color{black}{=~?} }\)

OpenStudy (solomonzelman):

multiply those together as i showed.

OpenStudy (noah690):

is the answer 4x^10-12x^5